OpenStudy (anonymous):

plz help me in ques no 1

OpenStudy (anonymous):

Q1 problem 2 ; Tr= -Rl*Cgs*ln((Vs-Voh)/Vs) Q1 problem 5: Tf= -Rth*Cgs*ln((Vol-Vth)/(Vs-Vth)) Q1 Problem 1: Vol= (Vs*Ron)/(Ron+r), Q1 Problem 3; Rl||Ron probl 4; Tau= Rth*Cgs, where Rth= Ron||Rl

OpenStudy (anonymous):

what is vth?

OpenStudy (anonymous):

how to get Vth..

OpenStudy (anonymous):

how to get Rth

OpenStudy (anonymous):

Q1p5?

OpenStudy (anonymous):

@axi what is vth how to calculate this, @afridi112